OpenStudy (anonymous):

Scores on an exam are normally distributed with a mean of 76 and a standard deviation of 10. In a group of 230 tests, how many students score below 66? about 93 students about 115 students about 37 students about 78 students

OpenStudy (amistre64):

how many standard deviations is 66 away from the mean?

OpenStudy (amistre64):

76 + x(sd) = 66 x(sd) = 66 - 76 x = (66-76)/(sd)

OpenStudy (amistre64):

then depending on accuaracy we could possibly use the empirical rule to determine the rest